OpenStudy (anonymous):

Bart bought a digital camera with a list price of $219 from an online store offering a 6 percent discount. He needs to pay $7.50 for shipping. What was Bart's total cost

OpenStudy (pinklion23):

im sorry but this question was already answered idk if it will help any but http://openstudy.com/study#/updates/4f2fdeb0e4b0571e9cbb6204

OpenStudy (texaschic101):

6 percent discount : 219 (.06) = 13.14 price with discount : 219 - 13.14 = 205.86 205.86 + s&h = total cost 205.86 + 7.50 = 213.36